Consider the following Data:

Year

Tea
(L per person)

Coffee
(L per person)

1994

42.4

95.85

1995

42.12

97.28

1996

47.61

87.62

1997

60.86

92.04

1998

55.58

99.21

1999

50.61

95.63

2000

49.89

97.42

2001

56.77

93.93

2002

62.53

95.67

2003

68.31

99.25

2004

69.88

101.31

2005

72.99

101.68

2006

71.36

104.02

2007

90.78

106.09

2008

74.7

105.8

2009

67.15

102.15

2010

67.03

101.15

2011

87.83

104.05

2012

93.4

102.7

2013

78.9

105.28

2014

111.32

106.3

2015

98.39

104.96

2016

105.25

103.57

By using the definition and discussing what is relevant to the situation, interpret each of the following for both the coffee and tea data. Also, compare each for coffee and tea. Be sure to include the relevant information (state the value of or, in the case of the distribution, include the graphs) with each component.

  1. Mean
  2. Median
  3. Modal Interval
  4. Range
  5. IQR
  6. Standard Deviation
  7. Distribution of histogram and box plot
  8. Slope of each linear model
  9. Y-intercept of Coffee vs. Tea
  10. Correlation coefficient for each linear model
  11. Relevant interpolations or extrapolations
  12. Correlation type for coffee and tea
Answer #1

Here I attach the R code with output

tea=c(42.4,42.12,47.61,60.86,55.58,50.61,49.89,56.77,62.53,68.31,69.88,72.99,71.36,90.78,74.7,67.15,67.03,87.83,93.4,78.9,111.32,98.39,105.25)
coffee=c(95.85,97.28,87.62,92.04,99.21,95.63,97.42,93.93,95.67,99.27,101.31,101.68,104.02,106.09,105.8,102.15,101.15,104.05,102.7,105.28,106.3,104.96,103.57)

year=c(1994:2016)  
mean(tea)
median(tea)
mfv1(tea)
range(tea)
mean(coffee)
median(coffee)
mfv1(coffee)

summary(tea)
summary(coffee)
IQRtea=83.36-56.18=27.18
IQRcoffee=104.00-96.56=7.44
sd(tea)
sd(coffee)

range(coffee)
hist(coffee)
hist(tea)
boxplot(coffee)
boxplot(tea)
plot(coffee,tea)

plot(year,coffee)
plot(year,tea)

m=lm(coffee~tea)
summary(m)
cor(coffee,tea)

The measure of central tendency are mean, median , mode(mfv-most frequent observation) etc..

Measure of spread is range

> mean (tea) [1] 70.68087 median (tea) 1] 68. 31 mfv1 (tea) [1] 42.12 range (tea) [1] 42.12 111.32 mean (coffee) [1] 100.1296

> summary (tea) Min. 1st Qu. Median Mean 3rd Qu 83. 36 111.30 мах. 42. 12 56. 18 68. 31 70. 68 > summary (coffee) Min. 1st Qu

range(tea)=111.32-42.12=69.2

range(coffee)=106.30-87.62=18.68

Histogram of coffee 85 90 95 100 105 110 coffee Frequency 9

Here the model interval length is 5 units.

Histogram of tea LO 80 40 60 100 120 tea Frequency 9

Here the model interval length is taken as 10 units.

Boxplot of coffee

From the plot it is clear that the distribution is negatively skewed and there is no outliers in the data.

95 000 06

Boxplot of tea

From the plot we can infer that the distribution is positively skewed and there is no outliers in the model.

00 08 09 O40

90 95 100 105 coffee C 08 09 O40 00 tea

1995 2000 2005 2010 2015 year coffee 06 105 000
1995 2000 2005 2010 2015 year 00 08 09 40 tea

The correlation between Coffee and tea is 0.769, which implies there is positive linear relatioship exist between the variable tea and coffee.

Call 1m (formula coffee tea) Residuals: Min 1Q Median 3Q Маx -8.002 -1.819 1.099 2.032 4.885 Coefficients: Estimate Std. Erro

The obtained model is:

coffee=86.3200+0.1954*tea